Skip to main content

AAVSOnet status, 2017-11-08

1 post / 0 new
HQA
HQA's picture
AAVSOnet status, 2017-11-08

All 5 active BSMs are fully operational.  Additional filters are being ordered for BSM_NH, now that it has a 10-position filter wheel.  (Everyone keep an eye out for SBIG CFW-10 filter wheels.  We need 3 more of them!)   Gary Walker is looking into alternatives for an enclosure for BSM_Hamren, as it is in a VERY moist climate.

TMO61 is off-line.  Jon Holzman is having communication problems with the MaxDome hardware for controlling the dome, and is investigating new serial-to-USB adapters.  We hope to get this running shortly.

OC61 is operating normally.

John Gross inadvertently restored a 2015 snapshot of the SRO system, and so lost recent plans.  We've been working to restore the plans based on a more recent dump of the Scheduler database (thanks, George!).  Since major changes needed to be made, John Gross is also updating to ACP V8 (Expert) at the same time.  It will take a few more days to iron out the operations.  The telescope is in good shape.  The weather pattern is now morphing into an west-to-east winter flow, which will mean fewer photometric nights for a while.

Arne

AAVSO 49 Bay State Rd. Cambridge, MA 02138 aavso@aavso.org 617-354-0484